# InfAlign: Inference-aware language model alignment

### v2

@@ 4.3 Finding better transformations for BoN and WoN
$\texttt{BoN}(\pi)({\bm{y}}\mid{\bm{x}})\propto\pi({\bm{y}}\mid{\bm{x}})\cdot\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}})^{N-1}.$
$\texttt{WoN}(\pi)({\bm{y}}\mid{\bm{x}})\propto\pi({\bm{y}}\mid{\bm{x}})\cdot(1-\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}}))^{N-1}.$
@@
For any transformation function $\Phi$, the solution $\pi^{*}_{\mathcal{R}_{\Phi},\beta}$ to the KL-regularized RL problem defined in Eq. 13 satisfies the followings: Let $F_{\Phi,\beta}(u)=\frac{\int_{0}^{u}e^{\Phi(u^{\prime})/\beta}{\rm d}u^{\prime}}{\int_{0}^{1}e^{\Phi(u^{\prime})/\beta}{\rm d}u^{\prime}}$. Then the following hold true:
  • The best-of-$N$ win rate over the base policy satisfies for any ${\bm{x}}$,
    $W^{\texttt{BoN}}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})=1-N\int_{0}^{1}F_{\Phi,\beta}(u)^{N}u^{N-1}{\rm d}u,$
  • The worst-of-$N$ win rate over the base policy satisfies for any ${\bm{x}}$,
    $W^{\texttt{WoN}}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})=N\int_{0}^{1}\left(1-F_{\Phi,\beta}(u)\right)^{N}(1-u)^{N-1}{\rm d}u,$
  • The KL divergence between $\pi^{*}_{r_{\Phi},\beta}$ and ${\pi_{\rm ref}}$ satisfies
    $D_{\text{KL}}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\|{\pi_{\rm ref}})=\frac{1}{\beta}\frac{\int_{0}^{1}\Phi(u)e^{\Phi(u)/\beta}{\rm d}u}{\int_{0}^{1}e^{\Phi(u)/\beta}{\rm d}u}-\log\left(\int_{0}^{1}e^{\Phi(u)/\beta}{\rm d}u\right).$ (14)
By varying $\beta$ in Eq. 13, we can obtain an alignment curve plotting the inference-time win rate and KL divergence deviation for different aligned policies. This allows us to compare the performance of different transformation functions $\Phi$.
In the rest of the section, we will consider different types of transformations, and analytically compute the alignment curves using Theorem 3 by varying $\beta$, i.e., the plot of $(D_{\text{KL}}(\pi^{*}_{r_{\Phi},\beta}\|{\pi_{\rm ref}}),W^{\rm T}_{r}(\pi^{*}_{r_{\Phi},\beta})\succ{\pi_{\rm ref}})$ for different $\beta$’s. The transformations we consider include optimal transformations for standard win rate, exponential functions, and optimization-based transformations, described below.
Optimal reward transformations for standard win rate.
The identity mapping proposed by Azar et al. (2023) and the logarithmic mapping as used by BoN distillation (Beirami et al., 2024; Yang et al., 2024; Amini et al., 2024; Sessa et al., 2024) are shown to be (almost) optimal for the standard win rate. We would like to understand whether they are still good candidates when inference-time procedures are considered.
Deriving an optimized reward transformation function.
For calibrated inference-time procedures like BoN, and WoN, due to Theorem 2, we have that the optimal reward transformation $\Phi$ is independent of the base policy. One can optimize for good $\Phi$’s using simple toy language models, and we have the following corollary.
@@
For any $\beta>0$, the $\Phi$ that achieves the optimal BoN win-rate vs. KL tradeoff must satisfy the following pair of equations:
$\Phi_{\texttt{BoN}}(u)=-N^{2}\int_{u}^{1}F_{\pi}(v)^{N-1}v^{N-1}{\rm d}v,$
@@
and for WoN, it satisfies that
$\Phi_{\texttt{WoN}}(u)=-N^{2}\int_{u}^{1}(1-v)^{N-1}(1-F_{\pi}(v))^{N-1}{\rm d}v,$
@@
Hence one can find a transformation function based on finding the fixed point of the coupled equations in Corollary 2 through iterative updates.
Figure 1: Best-of-$N$ and worst-of-$N$ win rate vs KL tradeoff curves for $N=2,4$ with different transformation functions.
Exponential tilting for reward transformation.
In addition to deriving the optimized transformation, motivated by the exponential tilting of loss functions (Li et al., 2021; 2023), we consider the following exponential transformation:
@@
where $\text{sign}(t)=1$ for $t\geq 0$ and $\text{sign}(t)=-1$ for $t<0.$ These exponential transformations are essentially helping to optimize different quantiles of the reward for different values of $t$ (Li et al., 2023). For a positive value of $t$, the exponential tilting transformation focuses on optimizing the high quantiles of the objective (calibrated reward) and indeed recovers the max function for a large positive $t$. Hence, we expect that positive values of $t$ help with BoN inference-time procedure. On the other hand, for a negative value of $t$, the transformation is akin to optimizing the lower quantiles of the objective (calibrated reward), which makes it a suitable transformation for the WoN inference-time procedure.
@@ Results.
In Fig. 1, we consider an ideal continuous language model and compare different alignment objectives on inference-time win rate vs KL divergence for three inference-time procedures: {standard, BoN, WoN }. The tradeoff curves are obtained by varying the strength of the regularizer, $\beta$. The different objectives correspond to different transformations of the calibrated reward, which include (1) identity mapping $\Phi=I(\cdot)$ (IPO); (2) logarithmic mapping $\Phi=\log$ (best-of-$N$ distillation); (3) exponential tilting $\Phi_{t}(\cdot)$ as defined in Eq. 15; (4) optimized transformation function by solving for fixed points in Corollary 2.
Next, we consider Best-of-$2$ and Best-of-$4$ win rate. Here, additionally we include bon_opt, which is obtained by deriving the fixed point of Corollary 2. As can be seen, the identity transformation is no longer optimal. The best tradeoffs are given by bon_opt. We also observe that $\exp(5x)$ and $\exp(10x)$ are almost as good as bon_opt for Best-of-$2$ and Best-of-$4,$ respectively. Moreover,the identity transformation and logarithmic transformation are sub-optimal in these cases, which shows that considering standard win rate as the only metric is not optimal when inference-time procedure is concerned. We also observe that the behavior of identity transformation and logarithmic transformation is different in that the identity transformation gives better tradeoffs.
@@
Finally, we consider Worst-of-$2$ and Worst-of-$4$ win rate. Again, it can be observed that bon_opt gives the best tradeoffs for this inference-time procedure. Here, $\exp(-5x)$ and $\exp(-10x)$ are almost as good for Worst-of-$2$ and Worst-of-$4,$ respectively. We also observe that identity transformation and logarithmic transformation are sub-optimal in these cases and the logarithmic transformation gives better tradeoffs for WoN compared to the identity transformation.
The above results demonstrate the importance of considering the inference-time procedure when performing alignment. We find that exponential transformation with different $t$’s are good for different inference-time procedures, which will be our focus in practical experiments. Next, we will examine whether a good transformation that we found on the idealized continuous language model generalizes in the wild to real-world scenarios. Before moving on to the experiments, we will have to offer a practical algorithm for solving the inference-time KL-regularized RL optimization problem, which is the subject of the next section.
5 CTRL: Calibrate-and-Transform Reinforcement Learning
In this section, we propose the CTRL method, which is our proposed solver for the inference-time win rate optimization problem. Recall from the previous section that the proposed solution could be decomposed into three stages: reward calibration and reward transformation followed by the standard KL-RL solver. In the previous section, we highlighted the reward transformation. In the rest of this section, we focus on approximate empirical calibration. Combining it with reward transformation yields the final CTRL algorithm, given in Algorithm 1.
Algorithm 1 CTRL Algorithm for BoN and WoN
1:Base policy ${\pi_{\rm ref}}$, (uncalibrated) reward model $r$, training prompts ${\bm{x}}\in\mathcal{X}$, number of offline rollouts per prompt $m$.
2:Compute approximate empirical calibration function $\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}$ using Algorithm 2 with $m$ offline rollouts per prompt.
3:Transform calibrated reward using exponential function per Eq. 15 with $t>0$ for BoN and $t<0$ for WoN.
4:Optimize KL-RL using calibrated and transformed reward per Eq. 6.
@@ 5 CTRL: Calibrate-and-Transform Reinforcement Learning
Consider empirical calibration, where we draw $K$ samples ${\bm{z}}_{1},{\bm{z}}_{2},...,{\bm{z}}_{K}$ from the reference model ${\pi_{\rm ref}}$ for each prompt ${\bm{x}}$ in the RL training data. We then sort the rewards to all the responses $\{r({\bm{x}},{\bm{z}}_{1}),r({\bm{x}},{\bm{z}}_{2}),...r({\bm{x}},{\bm{z}}_{K})\}$, and assign empirical calibrated reward scores during RLHF training for the prompt, response pair $({\bm{x}},{\bm{y}})$ as
$\widetilde{\mathcal{C}}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})=\frac{1}{K}\sum_{i=1,z_{i}\sim{\pi_{\rm ref}}}^{K}\mathbbm{1}[r({\bm{x}},{\bm{y}})\geq r({\bm{x}},{\bm{z}}_{i})].$ (16)
Ideally, as $K\rightarrow\infty$, the empirical calibrated reward would converge to the true calibrated reward, and may be used in the RL training objective through PPO (Schulman et al., 2017). However, this could be costly as the exact computation of this calibrated reward requires us to sample and store $K$ reward scores per-prompt and per roll-out in the KL-RL solvers.
Instead, we propose to approximate the calibration curve by scaling it with a step-wise function in the logarithmic domain. We do this by choosing $p$ anchor points $q_{1},q_{2},\ldots$, where at each of the quantiles $q_{i}\in(0,1)$ we achieve zero calibration error. The algorithm for the simpler case ($p=1$, median) is given in Algorithm 2. For larger values of $p$, the algorithm is given in Algorithm 3 (Appendix D).
Algorithm 2 Offline Approximate Calibration
@@
1:Base policy ${\pi_{\rm ref}}$, (uncalibrated) reward model $r$ with range $[0,1]$, training prompts ${\bm{x}}\in\mathcal{X}$, number of samples per prompt $m$.
2:for all ${\bm{x}}\in\mathcal{X}$ do
3:     Roll out $m$ i.i.d. samples from ${\pi_{\rm ref}}$ per training prompt ${\bm{x}}$: ${\bm{z}}_{j},\,j\in\{1...m\}$.
4:     $r_{\text{median}}({\bm{x}})\leftarrow\text{median}(r({\bm{x}},{\bm{z}}_{1}),\ldots,r({\bm{x}},{\bm{z}}_{m}))$.
5:end for
@@
6:Define the calibrated reward function $\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}:\mathcal{X}\times\mathcal{Y}\to[0,1]$ as follows:
$\log\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})=\log r({\bm{x}},{\bm{y}})\cdot\frac{\log 0.5}{\log r_{\text{median}}({\bm{x}})}.$
7:Output: Calibrated reward function $\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}$ that assigns a calibrated reward for all prompts in the training set.
@@ 6.1 Evaluation setup
To train the reward models, we use the Anthropic Helpfulness and Harmlessness datasets (Bai et al., 2022) which involve multi-turn dialogues between a human and a digital assistant. The preference datasets consist of two responses for one context, and a label for the human preference for the response. We use the train split of the two datasets (44K examples for helpfulness and 42K for harmlessness) to train the uncalibrated and calibrated reward models – separate reward models for each objective.
@@
The uncalibrated reward model is trained based on the Bradley-Terry pairwise objective (Raffel et al., 2020), and the calibration is done on the training-split of the RL training procedure by drawing samples from the reference model. The underlying model for both these rewards is the PaLM-2 S model (Anil et al., 2023). The base reference policy model is a PaLM-2 S model that is fine-tuned (SFT) on the Anthropic dialog preferred responses. We then train the aligned policy model through KL-regularized Reinforcement Learning (Schulman et al., 2017; Bai et al., 2022). We compare against uncalibrated (a model trained with KL-RL using PPO and no further processing), BoNBoN (Gui et al., 2024), Best-of-N (Nakano et al., 2022; Beirami et al., 2024), BonD (Sessa et al., 2024), and IPO (Azar et al., 2023) as baselines.
@@
As evaluating using ground truth rewards in a pointwise manner based on human annotations can be expensive, we follow Eisenstein et al. (2024); Mudgal et al. (2024) and perform automated evaluation using a larger PaLM-2 M model to compute true rewards.
@@
To measure improvement due to post-RL training, we report both the win rate and the BoN and WoN win rates, calculated after applying the respective inference-time procedures on the responses generated by the RL model against the base SFT model, along with the corresponding KL-divergence of the RL model with the SFT model. For each of the runs, we experiment with different KL-regularizer strengths ($\beta\in\{0.01,0.02,\ldots,0.09\}$) and obtain the Pareto-curve of the KL divergence vs {standard, BoN, WoN } win rate curves.

### v5

@@ 4.3 Finding reward transformations for BoN and WoN
$\texttt{BoN}_{\pi}({\bm{y}}\mid{\bm{x}})=N\cdot\pi({\bm{y}}\mid{\bm{x}})\cdot\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}})^{N-1}.$
$\texttt{WoN}_{\pi}({\bm{y}}\mid{\bm{x}})=N\cdot\pi({\bm{y}}\mid{\bm{x}})\cdot(1-\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}}))^{N-1}.$
@@
For any transformation function $\Phi$, the solution $\pi^{*}_{\mathcal{R}_{\Phi},\beta}$ to InfAlign-CTRL satisfies the followings: Let $F_{\Phi,\beta}(u)=\frac{\int_{0}^{u}e^{\Phi(u^{\prime})/\beta}{\rm d}u^{\prime}}{\int_{0}^{1}e^{\Phi(u^{\prime})/\beta}{\rm d}u^{\prime}}$.
  • For any ${\bm{x}},W^{\texttt{BoN}}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})=$
    $1-N\int_{0}^{1}F_{\Phi,\beta}(u)^{N}u^{N-1}{\rm d}u,\vskip-7.22743pt$ (8)
  • For any ${\bm{x}},W^{\texttt{WoN}}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})=$
    $N\int_{0}^{1}\left(1-F_{\Phi,\beta}(u)\right)^{N}(1-u)^{N-1}{\rm d}u,\vskip-7.22743pt$ (9)
  • $D_{\text{KL}}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\|{\pi_{\rm ref}})=$
    $\frac{1}{\beta}\frac{\int_{0}^{1}\Phi(u)e^{\Phi(u)/\beta}{\rm d}u}{\int_{0}^{1}e^{\Phi(u)/\beta}{\rm d}u}-\log\left(\int_{0}^{1}e^{\Phi(u)/\beta}{\rm d}u\right).$ (10)
The above theorem generalizes the win rate calculation in Gui et al. (2024, Lemma 5) to inference-time win rate. By varying $\beta$ in Theorem 3, we obtain an alignment curve plotting the inference-time win rate and KL divergence for different aligned policies. This allows us to compare the performance of different transformation functions.
In the rest of the section, we investigate different types of transformations, and analytically compute the alignment curves, i.e., the plot of $(D_{\text{KL}}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\|{\pi_{\rm ref}}),W^{T}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}})$ for different $\beta$’s. The transformations we consider include optimal transformations for standard win rate, exponential functions, and optimization-based transformations.
Optimal reward transformations for standard win rate. The identity mapping $\Phi(x)=x$ proposed by Azar et al. (2023) and the logarithmic mapping $\Phi(x)=\log x$ as used by BoN distillation (Beirami et al., 2024; Yang et al., 2024; Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024) are shown to be (almost) optimal for the standard win rate. We investigate whether these transformations are suited to inference-time procedures.
Deriving an optimized reward transformation function. Due to Theorem 2, one can optimize for good $\Phi$’s using simple toy language models, leading to the following corollary.
@@
For any $\beta>0$, the $\Phi$ that solves InfAlign with $\mathcal{T}=\texttt{BoN}$ must satisfy the following pair of equations:
$\Phi_{\texttt{BoN}}(u)=-N^{2}\int_{u}^{1}F(v)^{N-1}v^{N-1}{\rm d}v,$
@@
where $F(v)=\int_{0}^{v}f(v){\rm d}v$ is the CDF with density function $f$. and for $\mathcal{T}=\texttt{WoN}$, it satisfies that
$\Phi_{\texttt{WoN}}(u)=-N^{2}\int_{u}^{1}(1-v)^{N-1}(1-F(v))^{N-1}{\rm d}v,$
@@
Hence, we derive a transformation function by iteratively finding the fixed point of the coupled equations in Corollary 2. We call them bon_fp and won_fp.
Figure 2: Best-of-NN (left) and Worst-of-$N$ (right) win rate vs KL tradeoff curves for $N=4$ with different transformation functions.
Exponential tilting for reward transformation. In addition to deriving the optimized transformation, motivated by the exponential tilting of loss functions (Li et al., 2021, 2023), we consider the following exponential transformation:
@@
where $\text{sign}(t)=1$ for $t\geq 0$ and $\text{sign}(t)=-1$ for $t<0.$ These exponential transformations are essentially helping to optimize different quantiles of the reward for different values of $t$ (Li et al., 2023). For a positive value of $t$, the exponential tilting transformation focuses on optimizing the higher quantiles of the objective (calibrated reward) . On the other hand, for a negative value of $t$, the transformation is akin to optimizing the lower quantiles of the calibrated reward, which makes it a suitable transformation for the WoN inference-time procedure.
@@ Results.
We compare different reward transformations on inference-time win rate vs KL divergence for three inference-time procedures: {standard, BoN, WoN } with different $N$’s. The tradeoff curves are obtained by varying the strength of the regularizer, $\beta$. We present the result for $N=4$ in Fig. 2 and provide more results in Appendix F.
@@
For Best-of-$4$ win rate, the identity and $\log$ transformation, which are optimal for standard win rate, are suboptimal. The best tradeoffs are given by bon_fp. We also observe that $\exp(10x)$ are almost as good as bon_fp for Best-of-$4$. For Worst-of-$4$ win rate, it is observed that won_fp gives the best tradeoffs for this inference-time procedure. Here, $\exp(-5x)$ and $\exp(-10x)$ are almost as good for Worst-of-$2$ and Worst-of-$4,$ respectively. We also observe that identity transformation and $\log$ transformation are sub-optimal in these cases and the $\log$ transformation gives better tradeoffs for WoN compared to the identity transformation.
The above results show that considering standard win rate as the only metric is not sufficient when inference-time procedure is concerned, demonstrating the importance of inference-aware alignment. We find that exponential transformation with different $t$’s are good for BoN and WoN procedures, which will be our focus in practical experiments.
4.4 Practical implementation of InfAlign-CTRL
@@
When implementing InfAlign-CTRL in practice, two questions remain to be resolved: (1) How to obtain calibrated reward $\mathcal{C}_{r,{\pi_{\rm ref}}}$; (2) How to solve the RLHF problem. To obtain $\mathcal{C}_{r,{\pi_{\rm ref}}}$, we consider empirical calibration, where we draw $K$ samples ${\bm{z}}_{1},{\bm{z}}_{2},...,{\bm{z}}_{K}$ from the reference model $\pi_{\rm ref}$} for each prompt $\bm{x}$} in the RL training data. We then sort the rewards to all the responses $\{r({\bm{x}},{\bm{z}}_{1}),r({\bm{x}},{\bm{z}}_{2}),...,r({\bm{x}},{\bm{z}}_{K})\}$, and assign empirical calibrated reward scores during RLHF training for the prompt, response pair $({\bm{x}},{\bm{y}})$ as
$\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})=\frac{1}{K}\sum_{i=1,{\bm{z}}_{i}\sim{\pi_{\rm ref}}}^{K}w_{r}({\bm{y}},{\bm{z}}_{i}\mid{\bm{x}}).\vskip-21.68121pt$ (12)
The following lemma establishes the error of approximating $\mathcal{C}_{r,{\pi_{\rm ref}}}$ with empirical calibration. The proof follows from DKW inequality (Dvoretzky et al., 1956).
Lemma 6.
Given ${\bm{x}}\in{\bm{\mathcal{X}}}$, for all $\delta>0$, with probabiltiy at least $1-\delta$,
$\max_{{\bm{y}}\in{\bm{\mathcal{Y}}}}|\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})-\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})|\leq\sqrt{\frac{\log(2/\delta)}{2K}}.$
For solving RLHF, we use PPO (Schulman et al., 2017) as the optimization algorithm in this paper to demonstrate the effectiveness of reward transformation. We believe the method could benefit from other advancements of RLHF algorithms.
Algorithm 1 Implementation of InfAlign-CTRL
@@
0: Base policy $\pi_{\rm ref}$}, (uncalibrated) reward model $r$, set of training prompts $D\subset{\bm{\mathcal{X}}}$, number of offline rollouts per prompt $K$, transformation function $\Phi$.
1: Compute empirical calibrated reward $\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}$ using Eq. 12 with $K$ offline rollouts per ${\bm{x}}\in D$.
@@
2: Transform calibrated reward using function $\Phi$ to get $\mathcal{R}_{\Phi}=\Phi\circ\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}$
3: Optimize RLHF using calibrated and transformed reward per Eq. 4 using PPO.
@@ 5.1 Evaluation setup
We consider the following tasks: (1) Anthropic Helpfulness and Harmlessness datasets (Bai et al., 2022), which involve multi-turn dialogues between a human and a digital assistant. For training the reward models, the preference datasets consist of two responses for one context, and a label for the human preference for the response. We use the train split of the two datasets (44K examples for helpfulness and 42K for harmlessness) to train the uncalibrated and calibrated reward models – separate reward models for each objective. (2) Similarly, for the summarization quality task, we use Reddit posts from TL;DR dataset (Stiennon et al., 2020) and train uncalibrated and calibrated reward models on the train split.
@@
The uncalibrated reward model is trained based on the Bradley-Terry pairwise objective (Raffel et al., 2020), and the calibration is done on the training-split of the RL training procedure. The underlying model for both these rewards is the PaLM-2 S model (Anil et al., 2023). The base reference policy model is a PaLM-2 S model that is fine-tuned (SFT) on the preferred responses of the Anthropic dialog and Reddit summarization datasets. We use InfAlign-CTRL with exponential transformations Eq. 11 and PPO as discussed in Algorithm 1 to obtain the aligned policy. We set $K=100$ in our experiments, and analyze the additional computational overhead in the Appendix.
Baselines.
We compare against uncalibrated (a model trained to solve RLHF with using the uncalibrated reward model using PPO), BoNBoN (Gui et al., 2024), BoND (Sessa et al., 2024), and IPO (Azar et al., 2023) as baselines.
@@
As evaluating using ground truth rewards in a pointwise manner based on human annotations can be expensive, we follow (Eisenstein et al., 2024; Mudgal et al., 2024) and perform automated evaluation using a larger PaLM-2 M model to compute true rewards and report the win-rate based on these rewards for responses generated by the aligned and base policy models. We acknowledge that there is a gap between these model-generated preference and human preference, but note that prior work (Bai et al., 2022; Stiennon et al., 2020) have reported inter-human-rater agreement on the 3 tasks is often less than 77% and correspondingly the model-generated true rewards have accuracy of (77.7%, 77.0, and 76.4%) on the Anthropic Helpfulness, Harmlessness, and Reddit text summarization preference datasets, comparable to state-of-the-art in RewardBench leaderboard (Lambert et al., 2025).
@@
To measure improvement due to post-RL training, we report both the win rate and the BoN and WoN win rates, along with the corresponding KL-divergence of the RL model with the SFT model. For each of the runs, we experiment with different KL-regularizer strengths ($\beta\in\{0.01,\ldots,0.09\}$) and obtain the Pareto-curve of the KL divergence vs {standard, BoN, WoN } win rate curves [^5] We use win-rate as our main evaluation metric. We present raw reward comparison for some experiments in Fig. 9 in the appendix..